# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ VI-D The elliptical parameterized “low-rank” model
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\mathbf{0},\bm{\Sigma}_{0},g_{0})$ be a zero-mean, RES-distributed vector whose scatter matrix $\bm{\Sigma}_{0}$ is modeled as:
@@
  • $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{R}^{m\times p}$ is a full rank column matrix with $m>p$, where $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ collects the parameter of interest and characterizes $\mathbf{A}_{0}$ and where the function $\mathbf{A}(\bm{\gamma})$ is differentiable,
@@
with $r=p(p+1)/2+1$, while the infinite dimensional nuisance is the density generator $g_{0}\in\overline{\mathcal{G}}$.
@@
$\mathcal{P}_{\bm{\gamma},\bm{\xi},g}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g)=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}g(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,g\in\overline{\mathcal{G}}\right\},$ (121)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g_{0})=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}g_{0}(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\}.$ (122)
Then, the following result can be proved by a direct application of Condition (115) of Proposition VI.1 as shown in Appendix .10.
@@
For the low-rank scatter model (119), the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$.
@@ VII Applications to circular and noncircular CES distributions
So far in this paper, we have only dealt with cases of real observation vectors. However complex-valued observations, i.e. $\mathbf{x}\in\mathbb{C}^{m}$ are an integral part of many science and engineering problems, including those in communications, radar, biomedicine, geophysics, oceanography, electromagnetics, and optics, among others. The complex field does not only provide a convenient representation for the observations but also provides a natural way to capture their physical nature as well as the transformations they go through (see e.g. [28, 29]). In many studies it has often been (implicitly) assumed that complex random vectors are circular, i.e. with invariant distribution under rotation around a center $\bm{\mu}$, that is $(\mathbf{x}-\bm{\mu})=_{d}e^{j\theta}(\mathbf{x}-\bm{\mu}),\;\forall\theta\in\mathbb{R}$ (see e.g. in [6]). This assumption however discards the information conveyed by the relationship between real and imaginary parts of the observation vectors. Consequently, the noncircularity may be an important feature that characterizes observation in many practical scenarios. For this reason, in this section, we believe it is important to provide some evidence to ensure that all results obtained so far in the case of real observations remain entirely valid for Complex Elliptically Symmetric (CES)-distributed observations. Moreover, given their importance in well-known engineering applications, we devote particular attention to the “complex” version of the Propositions VI.2 and VI.3. Before moving on, it is important to note that, although there are many different notations in the literature, in the following sections we will use the notation introduced in [8].
@@ VII-A Real-complex representations
where ${\bf M}\triangleq\tiny{\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}{\bf I}&i{\bf I}\\ {\bf I}&-i{\bf I}\\ \end{array}\right)}$ is a unitary matrix [8]. So, by indicating as $g_{r}\in\overline{\mathcal{G}}$ (defined in (34)) the density generator of the $2m$-dimensional real vector $\overline{\bf x}$ and if its scatter matrix $\overline{\bm{\Sigma}}\in\mathcal{S}_{2m}^{\mathbb{R}}$ is positive definite, then the pdf (29) of $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{C}^{m}$ is generally rewritten in the following form:
@@
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},g_{c})=|{\bf\Sigma}|^{-1}g_{c}\left[({\bf x}-{\bm{\mu}})^{H}{\bf\Sigma}^{-1}({\bf x}-{\bm{\mu}})\right].$ (127)
Using this complex representation of even-dimensional RES distributions, the Stochastic Representation Theorem (31) can be extended to both circular and non circular CES distributions as discussed in details in [8, Sect. 1.3.2]. These “complex” stochastic representations in fact use the mutually independent, random variable $\mathcal{Q}_{c}\triangleq\frac{1}{2}\mathcal{Q}_{r}$ (where $\mathcal{Q}_{r}$ is the 2nd-order modular variate associated with $\overline{\bf x}$) and the random vector ${\bf u}_{c}\sim\mathcal{U}(S_{\mathbb{C}}^{m-1})$. Moreover, the definition of the shape matrix (37) can be straightforwardly extended while keeping definition of matrix scale function in (36). In particular for NC-CES distributions, the shape matrix
@@
    Take the results obtained in the previous sections and consider the specific case of a real observed vector of dimension $2m$, i.e. $\mathbb{R}^{2m}\ni\overline{\bf x}\sim RES_{2m}(\overline{\bm{\mu}},\overline{\bm{\Sigma}},g_{r})$,
@@
    Consider directly a complex observation vector, i.e. $\mathbb{C}^{m}\ni{\bf x}\sim\text{{NC-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},{\bm{\Omega}},g_{c})$ in the case of NC-CES distributions or $\mathbb{C}^{m}\ni{\bf x}\sim\text{{C-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},g_{c})$ in the case of C-CES distributions.
@@
Exstensive discussions and related examples about the above mentioned recipes can be found in [33, 34, 35, 36, 37, 24].
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=\beta(g_{c,0})\mathbf{J}[\widetilde{\bm{\mu}}_{0}]^{H}\widetilde{\bm{\Sigma}}_{0}^{-1}\mathbf{J}[\widetilde{\bm{\mu}}_{0}],$ (128)
for NC-CES distributions where $\beta(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m}$ with $\varphi_{c,0}(t)\triangleq\frac{-1}{g_{c,0}(t)}\frac{dg_{c,0}(t)}{dt}$ and $\mathbf{J}[\widetilde{\bm{\mu}}_{0}]]_{i,j}\triangleq\left.\frac{\partial[\widetilde{\bm{\mu}}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$. In the particular case of C-CES distributions, it is immediate to verify that the FIM in (128) reduces to
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=2\beta(g_{c,0}){\rm Re}\left\{\mathbf{J}[\bm{\mu}_{0}]^{H}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\right\}$ (129)
@@
In signal processing, there are many examples of complex-valued observations ${\bf x}\in\mathbb{C}^{m}$ where the location vector $\bm{\mu}$ includes the parameters of interest $\bm{\gamma}$, while the scatter matrix $\bm{\Sigma}$ gathers the nuisance parameters $\bm{\xi}$. This is the case where a deterministic signal of interest parameterized by $\bm{\gamma}$ is disturbed by a zero-mean C-CES distributed noise, whose density generator $g$ is unspecified. A classic exemple is given by the statistical model for time delay and Doppler estimation problems (see e.g. [39, 40]).
@@
Proposition VI.3 also extends to C and NC-CES distributions by considering the one-to-one mapping $\overline{\bm{\Sigma}}\mapsto\widetilde{\bf\Sigma}=2{\bf M}\overline{\bm{\Sigma}}{\bf M}^{H}$. More specifically for the C-CES distribution, (119) can be rewritten in the following form [41]:
@@
where $\bm{\Sigma}_{0}\in{\cal S}_{m}^{\mathbb{C}}$, $\bm{\Xi}_{0}\in{\cal S}_{p}^{\mathbb{C}}$ are two symmetric and positive definite matrices and $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{C}^{m\times p}$ is full column rank that collects the parameters of interest $\bm{\gamma}_{0}$ that characterize $\mathbf{A}_{0}$.
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(g_{c,0})}{\lambda_{0}}{\rm Re}\left\{{\bf J}[\mathrm{vec}({\bf A}_{0})]^{H}({\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{{\bf A}_{0}}){\bf J}[\mathrm{vec}({\bf A}_{0})]\right\},$ (131)
@@
${\bf J}[\mathrm{vec}({\bf A}_{0})]]_{i,j}\triangleq\left.\frac{\partial[\mathrm{vec}({\bf A}(\bm{\gamma}))]_{i}}{\partial\gamma_{j}}\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ and where here $\alpha(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}^{2}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m(m+1)}$. Note that the parametric efficient FIM (131) reduces to:
$\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(g_{c,0})}{\lambda_{0}}{\rm Re}\left\{({\bf D}_{0}^{H}{\bf\Pi}^{\bot}_{{\bf A}_{0}}{\bf D}_{0})\odot{\bf H}_{0}^{T}\right\}$ (134)
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{\alpha(g_{c,0})}{\lambda_{0}}{\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})]^{H}(\widetilde{\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{\widetilde{\mathbf{A}}_{r,0}}){\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})],$ (138)

### v3

@@ VI-D The elliptical parameterized “low-rank” model
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\mathbf{0},\bm{\Sigma}_{0},\bar{g}_{0})$ be a zero-mean, RES-distributed vector whose scatter matrix $\bm{\Sigma}_{0}$ is modeled as:
@@
  • $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{R}^{m\times p}$ is a full rank column matrix with $m>p$, where $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ collects the parameters of interest which characterize $\mathbf{A}_{0}$ and where the function $\mathbf{A}(\bm{\gamma})$ is differentiable,
@@
with $r=p(p+1)/2+1$, while the infinite dimensional nuisance is the density generator $\bar{g}_{0}\in\overline{\mathcal{G}}$.
@@
$\mathcal{P}_{\bm{\gamma},\bm{\xi},\bar{g}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},\bar{g})=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}\bar{g}(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,\bar{g}\in\overline{\mathcal{G}}\right\},$ (129)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},\bar{g}_{0})=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}\bar{g}_{0}(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\}.$ (130)
The following result can be proved by a direct application of Condition (117) of Proposition VI.1 as shown in Appendix .11.
@@
For the low-rank scatter model (127), the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{0})$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$.
@@ VII Applications to circular and noncircular CES distributions
So far in this paper, we have only dealt with cases of real observation vectors. However complex-valued observations, i.e. $\mathbf{x}\in\mathbb{C}^{m}$ are an integral part of many science and engineering problems, including those in communications, radar, biomedicine, geophysics, oceanography, electromagnetics, and optics, among others. The complex field does not only provide a convenient representation for the observations but also provides a natural way to capture their physical nature as well as the transformations they go through (see e.g. [38, 39]). In many studies it has often been (implicitly) assumed that complex random vectors are circular, i.e. with invariant distribution under rotation around a center $\bm{\mu}$, that is $(\mathbf{x}-\bm{\mu})=_{d}e^{j\theta}(\mathbf{x}-\bm{\mu}),\;\forall\theta\in\mathbb{R}$ (see e.g. in [7]). This assumption however discards the information conveyed by the relationship between real and imaginary parts of the observation vectors. Consequently, the noncircularity may be an important feature that characterizes observation in many practical scenarios where the data are non-stationary. For this reason, in this section, we believe it is important to provide some evidence to ensure that all results obtained so far in the case of real observations remain entirely valid for Complex Elliptically Symmetric (CES)-distributed observations. Moreover, given their importance in well-known engineering applications, we devote particular attention to the “complex” version of the Propositions VI.2 and VI.3. Before moving on, it is important to note that, although there are many different notations in the literature, in the following sections we will use the notation introduced in [9].
@@ VII-A Real-complex representations
where ${\bf M}\triangleq\tiny{\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}{\bf I}&i{\bf I}\\ {\bf I}&-i{\bf I}\\ \end{array}\right)}$ is a unitary matrix [9]. So, by indicating as $\bar{g}_{r}\in\overline{\mathcal{G}}$ (defined in (34)) the density generator of the $2m$-dimensional real vector $\overline{\bf x}$ and if its scatter matrix $\overline{\bm{\Sigma}}\in\mathcal{S}_{2m}^{\mathbb{R}}$ is positive definite, then the pdf (29) of $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{C}^{m}$ is generally rewritten in the following form:
@@
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},\bar{g}_{c})=|{\bf\Sigma}|^{-1}\bar{g}_{c}\left[({\bf x}-{\bm{\mu}})^{H}{\bf\Sigma}^{-1}({\bf x}-{\bm{\mu}})\right].$ (135)
Using this complex representation of even-dimensional RES distributions, the Stochastic Representation Theorem (31) can be extended to both circular and non-circular CES distributions as discussed in details in [9, Sect. 1.3.2]. These “complex” stochastic representations in fact use the mutually independent, random variable $\mathcal{Q}_{c}\triangleq\frac{1}{2}\mathcal{Q}_{r}$ (where $\mathcal{Q}_{r}$ is the 2nd-order modular variate associated with $\overline{\bf x}$) and the random vector ${\bf u}_{c}\sim\mathcal{U}(S_{\mathbb{C}}^{m-1})$. Moreover, the definition of the shape matrix (37) can be straightforwardly extended while keeping definition of matrix scale function in (36). In particular for NC-CES distributions, the shape matrix
@@
    Take the results obtained in the previous sections and consider the specific case of a real observed vector of dimension $2m$, i.e. $\mathbb{R}^{2m}\ni\overline{\bf x}\sim RES_{2m}(\overline{\bm{\mu}},\overline{\bm{\Sigma}},\bar{g}_{r})$,
@@
    Consider directly a complex observation vector, i.e. $\mathbb{C}^{m}\ni{\bf x}\sim\text{{NC-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},{\bm{\Omega}},\bar{g}_{c})$ in the case of NC-CES distributions or $\mathbb{C}^{m}\ni{\bf x}\sim\text{{C-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},\bar{g}_{c})$ in the case of C-CES distributions.
@@
Extensive discussions and related examples about the above mentioned recipes can be found in [43, 44, 45, 46, 47, 26].
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=\beta(\bar{g}_{c,0})\mathbf{J}[\widetilde{\bm{\mu}}_{0}]^{H}\widetilde{\bm{\Sigma}}_{0}^{-1}\mathbf{J}[\widetilde{\bm{\mu}}_{0}],$ (136)
for NC-CES distributions where $\beta(\bar{g}_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}\bar{\varphi}^{2}_{c,0}(\mathcal{Q}_{c})\}}{m}$ with $\bar{\varphi}_{c,0}(t)\triangleq\frac{-1}{\bar{g}_{c,0}(t)}\frac{d\bar{g}_{c,0}(t)}{dt}$ and $\mathbf{J}[\widetilde{\bm{\mu}}_{0}]]_{i,j}\triangleq\left.\frac{\partial[\widetilde{\bm{\mu}}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$. In the particular case of C-CES distributions, it is immediate to verify that the FIM in (136) reduces to
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=2\beta(\bar{g}_{c,0}){\rm Re}\left\{\mathbf{J}[\bm{\mu}_{0}]^{H}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\right\}$ (137)
@@
In signal processing, there are many examples of complex-valued observations ${\bf x}\in\mathbb{C}^{m}$ where the location vector $\bm{\mu}$ includes the parameters of interest $\bm{\gamma}$, while the scatter matrix $\bm{\Sigma}$ gathers the nuisance parameters $\bm{\xi}$. This is the case where a deterministic signal of interest parameterized by $\bm{\gamma}$ is disturbed by a zero-mean C-CES distributed noise, whose density generator $\bar{g}$ is unspecified. A classic example is given by the statistical model for time delay and Doppler estimation problems (see e.g. [49, 50]).
@@
Proposition VI.3 also extends to C- and NC-CES distributions by considering the one-to-one mapping $\overline{\bm{\Sigma}}\mapsto\widetilde{\bf\Sigma}=2{\bf M}\overline{\bm{\Sigma}}{\bf M}^{H}$. More specifically for the C-CES distribution, (127) can be rewritten in the following form [51]:
@@
where $\bm{\Sigma}_{0}\in{\cal S}_{m}^{\mathbb{C}}$, $\bm{\Xi}_{0}\in{\cal S}_{p}^{\mathbb{C}}$ are two Hermitian and positive definite matrices and $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{C}^{m\times p}$ is full column rank that collects the parameters of interest $\bm{\gamma}_{0}$ that characterize $\mathbf{A}_{0}$.
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(\bar{g}_{c,0})}{\lambda_{0}}{\rm Re}\left\{{\bf J}[\mathrm{vec}({\bf A}_{0})]^{H}({\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{{\bf A}_{0}}){\bf J}[\mathrm{vec}({\bf A}_{0})]\right\},$ (139)
@@
${\bf J}[\mathrm{vec}({\bf A}_{0})]]_{i,j}\triangleq\left.\frac{\partial[\mathrm{vec}({\bf A}(\bm{\gamma}))]_{i}}{\partial\gamma_{j}}\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ and where here $\alpha(\bar{g}_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}^{2}\bar{\varphi}^{2}_{c,0}(\mathcal{Q}_{c})\}}{m(m+1)}$. Note that the parametric efficient FIM (139) reduces to:
$\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{c,0})=\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(\bar{g}_{c,0})}{\lambda_{0}}{\rm Re}\left\{({\bf D}_{0}^{H}{\bf\Pi}^{\bot}_{{\bf A}_{0}}{\bf D}_{0})\odot{\bf H}_{0}^{T}\right\}$ (142)
@@
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{\alpha(\bar{g}_{c,0})}{\lambda_{0}}{\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})]^{H}(\widetilde{\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{\widetilde{\mathbf{A}}_{r,0}}){\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})],$ (146)